# Active Human Feedback Collection via Neural Contextual Dueling Bandits

### v1

@@ 3.4 Active Dueling Bandits with Regret Minimization
In many real-world applications, such as medical treatment design (Lai & Robbins, 1985; Bengs et al., 2021) and content moderation (Avadhanula et al., 2022), both actively selecting arms and minimizing regret is required. For instance, in personalized medical treatment, active learning is used to identify the most informative treatments to test, while cumulative regret minimization ensures the system continually adapts to deliver better patient outcomes.
@@
Corollary 1 (UCB).
(Verma et al., 2025, Theorem 2) Let $\lambda>\kappa_{\mu}$ and $w\geq\text{poly}(T,L,K,1/\kappa_{\mu},L_{\mu},1/\lambda_{0},1/\lambda,\log(1/\delta))$. Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using UCB-based arm selection strategy is upper bounded by
Corollary 2 (TS).
(Verma et al., 2025, Theorem 3) Let the conditions in 1. Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using TS-based arm selection strategy is upper bounded by
@@ 4 Experiments
To validate our theoretical results, we empirically evaluate the performance of our algorithms on different problem instances of synthetic datasets. Specifically, we use two commonly used synthetic functions adopted from existing works on neural bandits (Zhou et al., 2020; Zhang et al., 2021; Dai et al., 2023; Verma et al., 2025): $f(x)=10(x^{\top}\theta)^{2}$ (Square), $f(x)=2\sin(x^{\top}\theta)$ (Sine), and $f(x)=\cos(3x^{\top}\theta)$ (Cosine). All experiments are repeated 10 times, and we report the average worst suboptimality gap with 95% confidence intervals (depicted as vertical lines on each curve). Additional results are also provided in Appendix A due to space constraints.
Synthetic dataset.   Let $x_{t,a}$ be the context-arm feature vector for context $c_{t}$ and an arm $a$. For all $t\geq 1$, $x_{t,a}$ is sampled uniformly at random from $(-1,1)^{d}$. We keep the number of arms constant across all rounds, denoted by $K$. In our experiments, the binary preference feedback indicating whether $x_{t,1}$ preferred over $x_{t,2}$ (representing human preference feedback) is sampled from a Bernoulli distribution with parameter $\mu\left(f(x_{t,1})-f(x_{t,2})\right)$).
Reward function estimation.   We use a neural network with $2$ hidden layers with width $50$ to estimate the latent reward function, $\lambda=1.0$, $\delta=0.05$, $d=20$, $K=10$, and fixed value of $\nu_{T}=\nu=1.0$ in all our experiments (unless we specifically indicate $d$ and $K$). Note that we did not perform any hyperparameter search for Neural-ADB.
@@
(g) Suboptimality Gap
(h) MAE
(i) Average Regret
Figure 1: Performance comparison under different synthetic functions, i.e., Square function (first row), Sine function (second row), and Cosine function (third row).
Comparison with baselines.   We compare the worst suboptimality gap (defined in Eq. 1), MAE (average suboptimality gap), and average regret (defined in Section 3.4) against the different baselines of active contextual dueling bandits to evaluate the performance of UCB- and TS-variant of Neural-ADB. We use three baselines: Random, AE-Borda (Mehta et al., 2023), APO (Das et al., 2024), AE-DPO (Mehta et al., 2023) and the neural variants of AE-Borda and AE-DPO in which we use a Neural network to estimate the latent reward function. They are named AE-Borda (NN) and Neural-ADB (APO) respectively. Experimental results in Fig. 1 show that our approach – Neural-ADB (UCB) – outperforms other baselines in almost all synthetic functions (i.e., square function and sine function) in terms of the suboptimality gap. Moreover, our approach outperforms other baselines on all synthetic functions in MAE and average regret. We have included more comparisons of our approach with other baselines in other settings (e.g., different $d$ or $K$) in Appendix B.
@@ 5 Related Work
Finite-Armed Dueling Bandits.   Learning from preference feedback has been extensively studied in the bandit literature. In the finite-armed dueling bandits setting, the learner aims to find the best arm while only observing preference feedback for two selected arms (Yue & Joachims, 2009; 2011; Yue et al., 2012). To determine the best arm in dueling bandits, different criteria, such as the Borda winner, Condorcet winner, Copeland winner, or von Neumann winner, have been used while focusing on minimizing regret using only pairwise preference feedback (Zoghi et al., 2014b; Ailon et al., 2014; Zoghi et al., 2014a; Komiyama et al., 2015; Gajane et al., 2015; Saha & Gopalan, 2018; 2019a; 2019b; Verma et al., 2019; 2020a; 2020b; Zhu et al., 2023). For a comprehensive overview of algorithms for various dueling bandits settings, we refer readers to the survey by Bengs et al. (2021).
@@
Contextual Dueling Bandits.   Many real-life applications, such as online recommendations, content moderation, medical treatment design, prompt optimization, and aligning large language models, can be effectively modeled using contextual dueling bandits, where a learner observes a context (additional information before selecting a pair of arms) and then selects the arms based on that context and observes preference feedback for the selected arms. Since the number of context-arm pairs can be potentially large or even infinite, the mean latent reward of each context-arm is assumed to be parameterized by an unknown function of its features. Common assumptions include linear reward models (Saha, 2021; Saha & Krishnamurthy, 2022; Bengs et al., 2022; Di et al., 2023; Li et al., 2024) and non-linear models (Verma et al., 2025). For our setting, we adopt the neural contextual dueling bandit algorithms proposed by Verma et al. (2025) to construct confidence ellipsoids for the latent non-linear reward function. Note that Neural-ADB can incorporate alternative confidence ellipsoids by appropriately modifying 2. Furthermore, our work addresses an active learning problem and analyzes the convergence rate of the worst sub-optimality gap, whereas Verma et al. (2025) focus on a regret minimization setting and derive upper bounds on cumulative regret.
@@
Active contextual dueling bandits.   The work most closely related to ours is active contextual dueling bandit (Mehta et al., 2023; Das et al., 2024), which takes a principled approach to actively collecting preference datasets. However, two key differences exist between our work and existing research: the non-linear reward function and the arm-selection strategy. Existing studies typically assume a linear reward function, which may not be suitable for many real-world applications. Our work addresses this gap by extending the existing framework to incorporate non-linear reward functions in contextual dueling bandits. Additionally, existing approaches use different methods for selecting the pair of arms, leading to distinct arm-selection strategies compared to ours. As a result of these differences in both the arm-selection strategy and the non-linear reward function (which we estimate using a neural network), our analysis diverges significantly from that of prior work.
@@ 6 Conclusion
This paper studies the problem of active human preference feedback collection by modeling it as an active neural contextual dueling bandit problem. We propose Neural-ADB, a principled and practical algorithm designed for efficiently gathering human preference feedback in scenarios where the reward function is non-linear. Exploiting the neural contextual dueling bandit framework, Neural-ADB extends its applicability to a broad range of real-world applications, including online recommendation systems and LLM alignment. Our theoretical analysis demonstrates that the worst suboptimality gap of Neural-ADB decays at a sub-linear rate as the preference dataset grows. Finally, our experimental results further validate these theoretical findings. An interesting direction for future work is applying Neural-ADB to real-life applications such as LLM alignment (Mehta et al., 2023; Das et al., 2024). From a theoretical perspective, exploring the non-stationary setting presents another promising future direction.
@@ Appendix A Leftover Proofs
    $\lambda_{\min}(V_{T})\geq\lambda+\sum_{s=1}^{T}\lambda_{\min}\left(z_{s}z_{s}^{\top}\right)$.
@@
    $\lambda_{\max}\left(z_{s}z_{s}^{\top}\right)\leq\left\|z_{s}\right\|_{2}^{2}\ \implies\ \lambda_{\max}(V_{T})\leq\lambda+\sum_{s=1}^{T}\left\|z_{s}\right\|_{2}^{2}$.
  4. 4.
    $\lambda_{\max}(V_{T}^{-1})={1}/{\lambda_{\min}(V_{T})}$.
  5. 5.
    $\lambda_{\max}(V_{T})=-{\lambda_{\min}(-V_{T})}$.
  6. 6.
@@
  7. 7.
    For any vector $z\in\mathbb{R}^{d}$, the Mahalanobis norm of $z$ with respect to $V_{T}$ is $\left\|z\right\|_{V_{T}}=\sqrt{z^{\top}V_{T}z}\leq\left\|z\right\|_{2}\sqrt{\lambda_{\max}(V_{T})}$.
  8. 8.
@@
  9. 9.
    Let $z_{\max}=\arg\!\max_{z\in\mathbb{R}^{d}}\left\|z\right\|_{V_{T}}$. Then, $z_{\max}$ is the eigenvector of $V_{T}$ corresponding to its largest eigenvalue.
  10. 10.
  11. 11.
@@ A.1 Proof of 1
Using Property 11 from 1 with $Y_{T}-\mathbb{E}\left[Y_{T}\right]=\mathbb{E}\left[V_{T}\right]-V_{T}$, we have
@@
Using upper bound on $\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)$ from Property 6 in 1, we get
@@
Therefore, with probability at least $1-\delta$, $\lambda_{\min}(V_{T})\geq T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\nicefrac{{d}}{{\delta}}\right)}$. Using Property 10 from 1, we now use to prove our key result as follows:
@@ A.3 Proof of 2 and 3
[…]
@@
$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}}}\left\|\frac{\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))}{\sqrt{w}}\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left\|\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi^{\prime}(c,a^{\star})-\varphi^{\prime}(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
@@
Appendix B Experimental Details and Additional Results
@@ B.1 Experimental Details
Practical considerations.   Based on the neural tangent kernel (NTK) theory (Jacot et al., 2018), the initial gradient $g(x;\theta_{0})$ can be used as the original feature vector $x$ as $g(x;\theta_{0})$ effectively represents the random Fourier features of the NTK. To make our algorithm more practical, we use common practices in neural bandits (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025). Specifically, we replaced the theoretical regularization parameter $\frac{1}{2}w\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2}$ (where $w$ is the NN’s width) with the simpler $\lambda\left\|\theta\right\|^{2}_{2}$ in the loss function (defined in Eq. 2) that is used to train our NN. We retrain the neural network after every $20$ rounds for $50$ gradient steps across all experiments.
B.2 Additional Results
Next, we present the additional experiment results comparing the performance of our proposed algorithms across varying input dimensions and different numbers of arms. We then evaluate the performance of Neural-ADB (TS) under various input dimensions and numbers of arms.

### v2

@@ 3.4 Active Dueling Bandits with Regret Minimization
In many real-world applications, such as medical treatment design (Lai and Robbins, 1985; Bengs et al., 2021) and content moderation (Avadhanula et al., 2022), both actively selecting arms and minimizing regret is required. For instance, in personalized medical treatment, active learning is used to identify the most informative treatments to test, while cumulative regret minimization ensures the system continually adapts to deliver better patient outcomes. Such scenarios also arise in other fields, such as dynamic pricing and personalized education, enabling systems to make smarter decisions, reduce suboptimal choices, and optimize overall performance as they gather more valuable data.
@@
Corollary 1 (Regret Upper Bound).
(Verma et al., 2025, Theorem 2 and Theorem 3) Let $\lambda>\kappa_{\mu}$ and $w\geq\text{poly}(T,L,K,1/\kappa_{\mu},L_{\mu},1/\lambda_{0},1/\lambda,\log(1/\delta))$. Then, with a probability of at least $1-\delta$, the regret of Neural-ADB when using UCB- or TS-based arm selection strategy is upper bounded by
@@ 4 Experiments
To validate our theoretical results, we empirically evaluate the performance of our algorithms on different problem instances of synthetic datasets. Specifically, we use two commonly used synthetic functions adopted from existing works on neural bandits (Zhou et al., 2020; Zhang et al., 2021; Dai et al., 2023; Verma et al., 2025): $f(x)=10(x^{\top}\theta)^{2}$ (Square) and $f(x)=2\sin(x^{\top}\theta)$ (Sine). All experiments are repeated 10 times, and we report the average worst suboptimality gap with 95% confidence intervals (depicted as vertical lines on each curve).
@@
Figure 1: Performance comparison of Neural-ADB against different active dueling bandit algorithms on synthetic functions: Square function (top row) and Sine function (bottom row).
(a) Suboptimality Gap
(b) MAE
(c) Average Regret
(d) Suboptimality Gap
(e) MAE
(f) Average Regret
Figure 2: Performance of Neural-ADB (UCB) on the Square function, evaluated across varying input dimensions (top row) and numbers of arms (bottom row).
(a) Sub-Optimality Gap
(b) MAE
(c) Average Regret
(d) Sub-Optimality Gap
(e) MAE
(f) Average Regret
Figure 3: Performance of Neural-ADB (TS) on the Square function, evaluated across varying input dimensions (top row) and numbers of arms (bottom row).
Synthetic dataset.   We generate sample features for each context-arm pair in a $d$-dimensional space. Let $x_{t,a}$ be the context-arm feature vector for context $c_{t}$ and an arm $a$. For all $t\geq 1$, $x_{t,a}$ is sampled uniformly at random from $(-1,1)^{d}$. We keep the number of arms constant across all rounds, denoted by $K$. In our experiments, the binary preference feedback indicating whether $x_{t,1}$ preferred over $x_{t,2}$ (representing human preference feedback) is sampled from a Bernoulli distribution with parameter $\mu\left(f(x_{t,1})-f(x_{t,2})\right)$), where $f$ is either a Square or Sine function..
Reward function estimation.   We use a neural network with $2$ hidden layers with width $50$ to estimate the latent reward function, $\lambda=1.0$, $\delta=0.05$, $d=20$, $K=10$, $T=1000$, and fixed value of $\nu_{T}=\nu=1.0$ in all our experiments (unless we specifically indicate $d$ and $K$). Note that we did not perform any hyperparameter search for Neural-ADB, whose performance can be further improved by doing the hyperparameter search.
Comparison with baselines.   We compare the worst suboptimality gap (defined in Eq. 1), MAE (average suboptimality gap, i.e., $\sum_{t=1}^{T}\left[\max_{a\in\mathcal{A}}f(\varphi(c_{t},a))-f\left(\varphi(c_{t},\pi(c))\right)\right]/T$), and average regret (defined in Section 3.4) against the different baselines of active contextual dueling bandits to evaluate the performance of UCB- and TS-variant of Neural-ADB. We use three baselines: Random, AE-Borda (Mehta et al., 2023), AE-DPO (Mehta et al., 2023), APO (Das et al., 2024), and the neural variants of AE-Borda and APO in which we use a neural network to estimate the latent reward function. They are named AE-Borda (NN) and Neural-ADB (APO) respectively. Experimental results in Fig. 1 show that our algorithm, Neural-ADB (UCB), outperforms other baselines in almost all synthetic functions (i.e., square function and sine function) in terms of the suboptimality gap. Moreover, both UCB- and TS-variants of Neural-ADB also outperform other baselines on all synthetic functions in MAE and average regret. We have included more comparisons of our approach with other baselines in other settings (e.g., different $d$ or $K$) in Appendix B.
[…]
@@ 5 Related Work
In the following, we briefly review the relevant work, especially in neural contextual bandits and dueling bandits, to our problem setting.
Neural Contextual Bandits.   To model complex and non-linear reward functions, neural contextual bandits (Riquelme et al., 2018; Zhou et al., 2020; Zhang et al., 2021; Xu et al., 2022; Bae and Lee, 2025; Verma et al., 2025) use deep neural networks for reward function estimation. (Riquelme et al., 2018) employ multi-layer neural networks to learn arm embeddings and then use Thompson Sampling at the final layer for exploration. Zhou et al. (2020) propose the first neural contextual bandit algorithm with sub-linear regret guarantees, using a UCB exploration strategy. Building on this, Zhang et al. (2021) propose an algorithm with a TS exploration strategy. Ban et al. (2022) introduces an adaptive exploration strategy incorporating an auxiliary neural network to estimate the potential gain of the exploitation neural network, diverging from traditional UCB and TS exploration strategies. To reduce the computational overhead of using gradient-based features, Xu et al. (2022) only perform UCB-based exploration on the final layer of the neural network. More recent works (Bae and Lee, 2025; Verma et al., 2025) extend these techniques to handle neural contextual bandit settings with binary feedback (i.e., neural logistic bandits).
@@
Finite-Armed Dueling Bandits.   Learning from preference feedback has been extensively studied in the bandit literature. In the finite-armed dueling bandits setting, the learner aims to find the best arm while only observing preference feedback for two selected arms (Yue and Joachims, 2009, 2011; Yue et al., 2012). To determine the best arm in dueling bandits, different criteria, such as the Borda winner, Condorcet winner, Copeland winner, or von Neumann winner, have been used while focusing on minimizing regret using only pairwise preference feedback (Ailon et al., 2014; Zoghi et al., 2014b, a; Gajane et al., 2015; Komiyama et al., 2015; Saha and Gopalan, 2018, 2019a, 2019b; Verma et al., 2019, 2020a, 2020b; Zhu et al., 2023). For a comprehensive overview of algorithms for various dueling bandits settings, we refer readers to the survey by Bengs et al. (2021).
@@
Contextual Dueling Bandits.   Many real-life applications, such as online recommendations, content moderation, medical treatment design, prompt optimization, and aligning large language models, can be effectively modeled using contextual dueling bandits, where a learner observes a context (additional information before selecting a pair of arms) and then selects the arms based on that context and observes preference feedback for the selected arms. Since the number of context-arm pairs can be potentially large or even infinite, the mean latent reward of each context-arm is assumed to be parameterized by an unknown function of its features. Common assumptions include linear reward models (Saha, 2021; Bengs et al., 2022; Di et al., 2023; Saha and Krishnamurthy, 2022; Li et al., 2024) and non-linear models (Verma et al., 2025). For our setting, we adopt the neural contextual dueling bandit algorithms proposed in (Verma et al., 2025) to construct confidence ellipsoids for the latent non-linear reward function. Note that Neural-ADB can incorporate alternative confidence ellipsoids by appropriately modifying 2. Furthermore, our work addresses an active learning problem and analyzes the convergence rate of the worst sub-optimality gap, whereas (Verma et al., 2025) focus on a regret minimization setting and derive upper bounds on cumulative regret.
@@
Active contextual dueling bandits.   The work most closely related to ours is active contextual dueling bandit (Mehta et al., 2023; Das et al., 2024), which takes a principled approach to actively collecting preference datasets. However, two key differences exist between our work and existing research: the non-linear reward function and the arm selection strategy. Existing studies typically assume a linear reward function, which may not be suitable for many real-world applications. Our work addresses this gap by extending the existing framework to incorporate non-linear reward functions in contextual dueling bandits. Additionally, existing approaches use different methods for selecting the pair of arms, leading to distinct arm selection strategies compared to ours. As a result of these differences in both the arm selection strategy and the non-linear reward function (which we estimate using a neural network), our analysis diverges significantly from that of prior work.
@@ 6 Conclusion
This paper studies the problem of active human preference feedback collection by modeling it as an active neural contextual dueling bandit problem. We propose Neural-ADB, a principled and practical algorithm designed for efficiently gathering human preference feedback in scenarios where the reward function is non-linear. Exploiting the neural contextual dueling bandit framework, Neural-ADB extends its applicability to a broad range of real-world applications, including online recommendation systems and LLM alignment. Our theoretical analysis demonstrates that the worst suboptimality gap of Neural-ADB decays at a sub-linear rate as the preference dataset grows. Finally, our experimental results further validate these theoretical findings. An interesting direction for future work is applying Neural-ADB to real-life applications such as LLM alignment. From a theoretical perspective, exploring the non-stationary setting presents another promising future direction.
@@ Appendix A Leftover Proofs
    $\lambda_{\max}(V_{T})=-{\lambda_{\min}(-V_{T})}$.
@@
  4. 4.
  5. 5.
@@ A.1 Proof of 1
Using Property 1 in 1 with $Y_{T}-\mathbb{E}\left[Y_{T}\right]=\mathbb{E}\left[V_{T}\right]-V_{T}$, we have
@@
Using upper bound on $\lambda_{\min}\left(V_{T}-\mathbb{E}\left[V_{T}\right]\right)$ from Property 3 in 1, we get
@@
Therefore, with probability at least $1-\delta$, $\lambda_{\min}(V_{T})\geq T\lambda_{\min}(\Sigma_{\max})-\sqrt{8\sum_{s=1}^{T}C_{s}\log\left(\nicefrac{{d}}{{\delta}}\right)}$. Using Property 4 in 1, we now use to prove our key result as follows:
@@ A.2 Proof of 1 and 2
See 2
Proof.
Recall that we are using the arm-selection strategies proposed in (Verma et al., 2025). Since their confidence bounds hold for any adapted sequence of contexts, the proof of the first part follows directly from Theorem 1 in (Verma et al., 2025), while the second part follows from Lemma 10 together with Eq. (27) of (Verma et al., 2025). ∎
Remark 1.
We adopt the arm selection strategies from the existing neural dueling bandit algorithms in (Verma et al., 2025), which assume $\tilde{d}=o(T)$. In some cases, $\tilde{d}=\Omega(T)$ (Ban et al., 2022; Deb et al., 2024), which may result in a constant convergence rate. However, our objective is to demonstrate the use of the neural network for estimating non-linear reward functions in active contextual dueling bandits. Since neural dueling bandit algorithms primarily influence the arm selection strategy, we can incorporate any variants of these algorithms by making appropriate modifications to 2.
@@ A.3 Proof of 2 and 3
$=\nu_{T}\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi(c,a^{\star})-\varphi(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
@@
$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}}}\left\|\frac{\varphi(c,a^{\star})-\varphi(c,\pi(c))}{\sqrt{w}}\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\max_{c\in\mathcal{C}}\left(\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\left\|\varphi(c,a^{\star})-\varphi(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
$=\nu_{T}\log\left(KT^{2}\right)\sqrt{\frac{\lambda}{\kappa_{\mu}w}}\max_{c\in\mathcal{C}}\left(\left\|\varphi(c,a^{\star})-\varphi(c,\pi(c))\right\|_{V_{T}^{-1}}\right)+2\varepsilon^{\prime}_{w,T}$
@@
Appendix B Additional Experimental Details and Results
@@ B.1 Experimental Details
Practical considerations.   Based on the neural tangent kernel (NTK) theory (Jacot et al., 2018), the initial gradient $g(x;\theta_{0})$ can be used as serve as a surrogate for the original feature vector $x$ as $g(x;\theta_{0})$ effectively represents the random Fourier features of the NTK. To make our algorithm more practical, we use common practices in neural bandits (Zhou et al., 2020; Zhang et al., 2021; Verma et al., 2025). Specifically, we replaced the theoretical regularization parameter $\frac{1}{2}w\lambda\left\|\theta-\theta_{0}\right\|^{2}_{2}$ (where $w$ is the NN’s width) with the simpler $\lambda\left\|\theta\right\|^{2}_{2}$ in the loss function (defined in Eq. 2) that is used to train our NN. We retrain the neural network after every $20$ rounds for $50$ gradient steps across all experiments.
B.2 Additional Experimental Results
Next, we present the additional experiment results comparing the performance of Neural-ADB varying input dimension $d$ (Fig. 4) and different numbers of arms $K$ (Fig. 5).